# On the Eligibility of LLMs for Counterfactual Reasoning: A Decompositional Study

### v1

@@ 4.1 LLM Performance on Decompositional Tasks (RQ1)
Furthermore, the results reveal a similar trend in modality influence as previously observed in Table 2, where complex, non-text modalities hinder effective reasoning. For example, in the code execution task (HumanEval-Exe), GPT-4o achieves only $<$0.65 F1 scores, which is notably lower than its performance on text-based tasks such as CLOMO.
Table 3: LLM performance (F1) in reasoning the counterfactual mediator ($M^{\prime}$) and outcome ($Y^{\prime}$).
Dataset GPT-4o Qwen-VL Llama-3.2-11B Gemini-Pro DeepSeek-VL
$M^{\prime}$ $Y^{\prime}$ $M^{\prime}$ $Y^{\prime}$ $M^{\prime}$ $Y^{\prime}$ $M^{\prime}$ $Y^{\prime}$ $M^{\prime}$ $Y^{\prime}$
CRASS 85.7±4.8 80.3±6.9 80.5±6.7 73.9±8.5 68.4±10.3 61.6±12.8 81.7±6.3 75.2±8.1 82.9±5.5 77.1±7.4
CLOMO 83.6±5.2 77.9±7.3 77.8±7.2 71.6±9.0 65.8±10.9 59.3±13.4 79.3±6.7 72.8±8.6 80.5±6.0 74.3±7.9
RNN-Topo 81.3±5.7 75.8±7.8 75.6±7.8 69.4±9.5 63.7±11.5 57.2±14.1 77.1±7.2 70.6±9.1 78.3±6.5 72.0±8.4
CVQA-Bool 71.8±9.3 65.2±11.4 65.4±11.2 58.6±13.3 53.1±14.8 46.5±17.2 63.2±12.1 56.8±14.0 66.7±10.4 59.8±12.5
CVQA-Count 68.4±10.1 61.9±12.2 62.2±12.0 55.7±14.1 50.3±15.6 43.8±18.0 60.1±12.9 53.8±14.9 63.5±11.2 56.9±13.3
COCO 66.2±10.7 59.5±12.9 60.1±12.7 53.4±14.8 48.0±16.3 41.5±18.7 57.8±13.6 51.5±15.6 61.3±11.9 54.6±14.0
Arithmetic 77.1±7.6 71.3±9.7 69.8±9.5 63.2±11.6 57.8±13.2 51.2±15.8 72.1±9.0 65.4±11.1 74.0±8.3 67.5±10.4
MalAlgoQA 74.8±8.2 68.9±10.3 67.5±10.1 60.9±12.2 55.9±13.8 49.3±16.4 69.6±9.7 62.9±11.8 71.8±8.9 65.1±11.0
HumanEval-Exe 64.4±11.3 57.6±13.5 58.2±13.3 51.5±15.4 46.1±17.0 39.6±19.4 55.8±14.2 49.4±16.2 59.4±12.5 52.7±14.7
Open-Critic 62.3±11.9 55.5±14.1 56.0±13.9 49.4±16.0 44.2±17.7 37.8±20.1 53.7±14.8 47.3±16.9 57.2±13.1 50.6±15.3
Code-Preference 63.5±11.6 56.7±13.8 57.1±13.6 50.4±15.7 45.1±17.3 38.7±19.8 54.7±14.5 48.3±16.5 58.3±12.8 51.6±15.0
Insights from counterfactual reasoning. Regardless of whether the setting is factual or counterfactual, the primary challenge lies not in identifying causal variables and performing causal reasoning. In particular, the complex input modality and the implicit nature of mediation hinder effective reasoning through causal pathways.
@@ 4.2 Auxiliary Techniques to Improve Counterfactual Reasoning (RQ2)
Dataset GPT-4o Llama-3.2-11B Gemini-Pro
$X$ $Z$ $Y$ $X$ $Z$ $Y$ $X$ $Z$ $Y$
CRASS $+$5.7 $+$6.3 $+$5.2 $+$16.5 $+$12.2 $+$12.6 $+$6.6 $+$5.3 $+$7.1
CLOMO $+$4.4 $+$4.8 $+$7.6 $+$22.6 $+$15.2 $+$18.3 $+$6.2 $+$9.4 $+$15.5
CVQA-Count $+$17.3 $+$15.5 $+$17.8 $+$33.6 $+$27.5 $+$25.6 $+$18.1 $+$15.7 $+$19.3
COCO $+$5.2 $+$2.8 $+$4.1 $+$7.8 $+$6.4 $+$10.3 $+$6.9 $+$3.8 $+$5.2
MalAlgoQA $+$3.9 $+$5.6 $+$2.3 $+$13.7 $+$10.9 $+$7.2 $+$12.4 $+$8.5 $+$8.3
Open-Critic $+$12.5 $+$14.2 $+$6.7 $+$18.9 $+$5.8 $+$1.7 $+$7.3 $+$6.1 $+$9.8
@@
4.2.1 Tool-Augmented Execution in Explicit Variable Identification Across Modalities
Settings. To enhance LLM performance in identifying explicit variables ($X,Z,Y$) across different modalities, we adopt a tool-augmented approach, where the LLM dynamically calls additional specialized tools to assist in entity identification, identical to a named entity recognition (NER) paradigm. We leverage several pretrained models from tailored to the multimodality of datasets:
@@ 4.2.1 Tool-Augmented Execution in Explicit Variable Identification Across Modalities
Experimental Results. We randomly select three representative LLMs and conduct experiments across multiple datasets. As shown in Table 4, tool-augmented execution consistently improves LLM performance in identifying explicit causal variables ($X,Z,Y$) across all modalities. For example, by leveraging GraphCodeBERT to parse code structures and forward the results to Llama, which gains a clearer understanding of programming logic and achieves an F1 improvement up to 0.189. Similarly, in the vision modality, the object detector grounding-dino-base assists by generating a set of candidate visual objects, which GPT-4o can then contextualize and compose into a coherent factual variable. For instance, detected objects like “Woman,” “Knife,” and “Apple” can be effectively integrated by GPT-4o into the causal expression: “A woman cuts an apple with a knife.”
@@ 4.2.2 Advanced Elicitation Strategies for Reasoning over Implicit Variables
Settings. To enhance LLM reasoning over implicit variables, particularly factual $M$ and counterfactual $M^{\prime},Y^{\prime}$, we implement advanced elicitation strategies that guide the model through more structured reasoning. Specifically, we apply Chain-of-Thought (CoT) wei2022chain , Chain-of-Thought with Self-Consistency (CoT-SC) wang2022self , and Tree-of-Thought (ToT) 10.5555/3666122.3666639 :
@@
In implementation, both CoT-SC and ToT are executed with $k$=5 sampled reasoning paths. ToT further evaluates candidate outputs by scoring their textual similarity using BERTScore bert-score , specifically assessing how well the intermediate results align with the original task statement.
Dataset Elicitation GPT-4o Llama-3.2-11B Gemini-Pro
$M$ $M^{\prime}$ $Y^{\prime}$ $M$ $M^{\prime}$ $Y^{\prime}$ $M$ $M^{\prime}$ $Y^{\prime}$
CoT $+$5.3 $+$3.7 $+$2.9 $+$8.6 $+$6.2 $+$4.5 $+$6.8 $+$4.5 $+$3.2
CRASS CoT-SC $+$5.1 $+$4.7 $+$5.3 $+$11.3 $+$8.9 $+$6.4 $+$9.4 $+$6.8 $+$5.0
ToT $+$6.5 $+$4.9 $+$3.8 $+$9.7 $+$8.1 $+$5.2 $+$8.2 $+$5.9 $+$4.3
CoT $+$6.1 $+$4.6 $+$3.3 $+$8.9 $+$7.5 $+$5.3 $+$7.6 $+$5.3 $+$3.9
CLOMO CoT-SC $+$6.7 $+$4.9 $+$3.9 $+$10.6 $+$8.3 $+$6.0 $+$8.9 $+$6.5 $+$4.7
ToT $+$9.8 $+$3.5 $+$2.6 $+$7.9 $+$5.8 $+$4.2 $+$12.4 $+$4.2 $+$3.0
CoT $+$5.4 $+$4.1 $+$3.2 $+$8.7 $+$6.8 $+$4.9 $+$7.3 $+$5.4 $+$3.8
CVQA-Count CoT-SC $+$3.8 $+$3.9 $+$2.1 $+$6.5 $+$7.7 $+$6.3 $+$12.2 $+$7.6 $+$4.5
ToT $+$4.6 $+$3.5 $+$2.8 $+$7.1 $+$5.9 $+$4.2 $+$6.1 $+$4.5 $+$3.1
CoT $+$3.5 $+$2.6 $+$1.9 $+$6.2 $+$4.3 $+$3.0 $+$4.9 $+$3.4 $+$2.3
COCO CoT-SC $+$5.1 $+$3.8 $+$2.9 $+$8.3 $+$6.4 $+$4.5 $+$6.9 $+$5.1 $+$3.6
ToT $+$4.2 $+$3.2 $+$2.4 $+$7.4 $+$5.8 $+$3.8 $+$5.6 $+$4.3 $+$3.0
CoT $+$4.3 $+$3.2 $+$2.4 $+$7.2 $+$5.2 $+$3.6 $+$5.8 $+$4.0 $+$2.7
MalAlgoQA CoT-SC $+$5.2 $+$4.0 $+$3.1 $+$8.3 $+$6.6 $+$4.7 $+$7.2 $+$5.2 $+$3.7
ToT $+$5.9 $+$4.5 $+$3.5 $+$9.5 $+$7.6 $+$5.4 $+$8.1 $+$6.0 $+$4.3
CoT $+$3.2 $+$2.4 $+$1.7 $+$5.8 $+$4.0 $+$2.8 $+$4.5 $+$3.1 $+$2.1
Open-Critic CoT-SC $+$3.9 $+$3.0 $+$2.3 $+$6.5 $+$4.8 $+$3.5 $+$5.0 $+$3.5 $+$2.7
ToT $+$4.7 $+$3.5 $+$2.6 $+$7.7 $+$6.0 $+$4.2 $+$6.3 $+$4.7 $+$3.4
@@ Appendix A Complementary Information of Causality and Causal Graph
CRASS Prompt.
# Role: Counterfactual Reasoning Analyst
Your task is to identify causal variables and generate causal graphs with edges. All outputs must follow strict JSON formatting for automated saving.
## Context Format
**Premise:**
{premise}
**Questionized Counterfactual Conditional (QCC):**
{qcc}
## Task Instructions
1. **Variable Identification**
- Extract and classify variables into:
- Exposure (X)‘: Direct intervention (verb phrase, e.g., "act of opening").
- Outcome (Y)‘: Final result (state phrase, e.g., "chest opened").
- Covariate (Z)‘: Pre-existing conditions influencing X/Y.
- Mediator (M)‘: Mechanism linking X-Y (physical/logical process).
2. **Causal Graph Construction**
- Build factual/counterfactual edges following these strict rules:
**Factual Graph Requirements:**
$\rightarrow$" Must include ALL these relationships:
1. Z - X (Covariate influences Exposure)
2. Z - M (Covariate influences Mediator)
3. Z - Y (Covariate influences Outcome)
4. X - M (Exposure affects Mediator)
5. M - Y (Mediator affects Outcome)
6. X - Y (Direct exposure effect)
**Counterfactual Graph Requirements:**
- Must include ALL these relationships:
1. Z - M (Same covariate influence on adjusted Mediator)
2. Z - Y (Same covariate influence on new Outcome)
3. X - M (Modified exposure affects adjusted Mediator)
4. M - Y (Adjusted mediator affects new Outcome)
5. X - Y (Direct modified exposure effect)
- Use prime notation (M, Y’) for counterfactual variables
- Maintain Zs original relationships
@@ Appendix B Complementary Experiment
This section presents additional experimental results that complement the main evaluation in the body of the paper. These supplementary findings, together with what has been presented in previous sections, offer comprehensive insights into LLMs capabilities across different decompositional tasks.
Appendix C Experimental Setting

### v2

@@ 4.1 LLM Performance on Decompositional Tasks (RQ1)
Insights from counterfactual reasoning. Regardless of whether the setting is factual or counterfactual, the primary challenge lies in identifying causal variables and performing causal reasoning. In particular, the complex input modality and the implicit nature of mediation hinder effective reasoning through causal pathways.
@@ 4.2 Auxiliary Techniques to Improve Counterfactual Reasoning (RQ2)
Dataset GPT-o4 Qwen3 Llama4-S Gemini2.5
$X$ $Z$ $Y$ $X$ $Z$ $Y$ $X$ $Z$ $Y$ $X$ $Z$ $Y$
CRASS ++6.0 ++6.6 ++5.5 ++10.8 ++9.4 ++9.7 ++15.2 ++11.5 ++11.9 ++6.6 ++5.3 ++7.1
CLOMO ++4.7 ++5.1 ++7.9 ++12.7 ++11.3 ++12.1 ++21.0 ++14.4 ++16.8 ++6.2 ++9.4 ++15.5
CVQA-Count ++17.7 ++15.9 ++18.2 ++21.9 ++21.4 ++22.7 ++32.0 ++26.1 ++24.1 ++18.1 ++15.7 ++19.3
COCO ++5.5 ++3.1 ++4.4 ++8.9 ++7.6 ++8.2 ++7.2 ++6.0 ++9.7 ++6.9 ++3.8 ++5.2
MalAlgoQA ++4.2 ++5.9 ++2.6 ++9.6 ++8.1 ++6.8 ++12.5 ++10.2 ++6.5 ++12.4 ++8.5 ++8.3
Open-Critic ++12.8 ++14.5 ++7.0 ++15.4 ++10.9 ++8.3 ++17.4 ++5.1 ++1.2 ++7.3 ++6.1 ++9.8
@@
4.2.1 Tool-Augmented Execution in Explicit Variable Identification
Settings. To enhance LLM performance in identifying explicit variables ($X,Z,Y$) across different modalities, we adopt a tool-augmented approach, where the LLM dynamically calls additional specialized tools to assist in entity identification, identical to a named entity recognition (NER) paradigm. We leverage several pretrained models tailored to the multimodality of datasets: (1) Text-based NER: We use bert-base-NER (Devlin et al., 2018; Tjong Kim Sang & De Meulder, 2003) to identify candidate entities in text-based data, including mathematical symbols.(2) Vision-based NER: We employ grounding-dino-base (Liu et al., 2023) to detect all relevant objects in images and generate focused regions by masking out irrelevant backgrounds. (3) Code analysis: We adopt GraphCodeBERT (Guo et al., 2020) to extract functions, variables, and control structures for programming tasks.
@@ 4.2.1 Tool-Augmented Execution in Explicit Variable Identification
Experimental Results. We randomly select three representative LLMs and conduct experiments across multiple datasets. As shown in Table 4, tool-augmented execution consistently improves LLM performance in identifying explicit causal variables ($X,Z,Y$) across all modalities. For example, by leveraging GraphCodeBERT to parse code structures and forward the results to Llama, which gains a clearer understanding of programming logic and achieves an F1 improvement up to 0.189. Similarly, in the vision modality, the object detector grounding-dino-base assists by generating a set of candidate visual objects, which GPT-o4 can then contextualize and compose into a coherent factual variable. For instance, detected objects like “Woman,” “Knife,” and “Apple” can be effectively integrated by GPT-o4 into the causal expression: “A woman cuts an apple with a knife.”
@@ 4.2.2 Advanced Elicitation Strategies for Reasoning over Implicit Variables
Dataset Elicitation GPT-o4 Qwen3 Llama4-S Gemini2.5
$M$ $M^{\prime}$ $Y^{\prime}$ $M$ $M^{\prime}$ $Y^{\prime}$ $M$ $M^{\prime}$ $Y^{\prime}$ $M$ $M^{\prime}$ $Y^{\prime}$
CoT ++5.6 ++4.0 ++3.1 ++7.4 ++5.6 ++4.1 ++7.9 ++5.8 ++4.2 ++6.8 ++4.5 ++3.2
CRASS CoT-SC ++5.3 ++5.0 ++5.5 ++8.6 ++7.3 ++6.0 ++10.5 ++8.2 ++5.7 ++9.4 ++6.8 ++5.0
ToT ++6.8 ++5.2 ++4.1 ++8.9 ++6.7 ++5.0 ++8.9 ++7.4 ++4.7 ++8.2 ++5.9 ++4.3
CoT ++6.4 ++4.9 ++3.6 ++8.3 ++6.5 ++4.8 ++8.2 ++6.9 ++5.0 ++7.6 ++5.3 ++3.9
CLOMO CoT-SC ++7.0 ++5.2 ++4.2 ++9.6 ++7.6 ++5.4 ++9.9 ++7.7 ++5.5 ++8.9 ++6.5 ++4.7
ToT ++10.1 ++3.8 ++2.9 ++8.7 ++5.2 ++3.5 ++7.2 ++5.1 ++3.9 ++12.4 ++4.2 ++3.0
CoT ++5.7 ++4.4 ++3.5 ++7.9 ++6.1 ++4.4 ++8.2 ++6.4 ++4.5 ++7.3 ++5.4 ++3.8
CVQA-Count CoT-SC ++4.1 ++4.2 ++2.4 ++7.2 ++6.6 ++4.8 ++5.9 ++6.8 ++5.7 ++12.2 ++7.6 ++4.5
ToT ++4.9 ++3.8 ++3.1 ++6.9 ++5.3 ++4.0 ++6.5 ++5.2 ++4.0 ++6.1 ++4.5 ++3.1
CoT ++3.8 ++2.9 ++2.1 ++6.0 ++4.6 ++3.2 ++5.4 ++4.0 ++2.6 ++4.9 ++3.4 ++2.3
COCO CoT-SC ++5.4 ++4.1 ++3.1 ++7.2 ++5.7 ++4.3 ++7.6 ++5.9 ++3.8 ++6.9 ++5.1 ++3.6
ToT ++4.5 ++3.5 ++2.7 ++6.8 ++5.1 ++3.6 ++6.2 ++5.0 ++3.5 ++5.6 ++4.3 ++3.0
CoT ++4.6 ++3.5 ++2.7 ++7.0 ++5.5 ++4.0 ++6.6 ++5.0 ++3.3 ++5.8 ++4.0 ++2.7
MalAlgoQA CoT-SC ++5.5 ++4.3 ++3.4 ++7.7 ++6.1 ++4.6 ++7.9 ++6.3 ++4.1 ++7.2 ++5.2 ++3.7
ToT ++6.2 ++4.8 ++3.8 ++8.5 ++6.7 ++5.0 ++8.6 ++7.1 ++4.6 ++8.1 ++6.0 ++4.3
CoT ++3.5 ++2.7 ++2.0 ++5.3 ++4.2 ++3.0 ++5.0 ++3.7 ++2.5 ++4.5 ++3.1 ++2.1
Open-Critic CoT-SC ++4.2 ++3.3 ++2.6 ++6.1 ++4.9 ++3.6 ++5.8 ++4.5 ++3.1 ++5.0 ++3.5 ++2.7
ToT ++5.0 ++3.8 ++2.9 ++6.7 ++5.2 ++3.8 ++6.1 ++5.3 ++3.6 ++6.3 ++4.7 ++3.4
Settings. To enhance LLM reasoning over implicit variables, particularly factual $M$ and counterfactual $M^{\prime},Y^{\prime}$, we implement advanced elicitation strategies that guide the model through more structured reasoning. Specifically, we apply Chain-of-Thought (CoT) (Wei et al., 2022), CoT with Self-Consistency (CoT-SC) (Wang et al., 2022), and Tree-of-Thought (ToT) (Yao et al., 2023):
@@
In implementation, both CoT-SC and ToT are executed with kk=5 sampled reasoning paths. ToT further evaluates candidate outputs by scoring their textual similarity using BERTScore (Zhang et al., 2019), specifically assessing how well the intermediate results align with the original task statement.
@@
4.2.3 Overall Performance Improvement
Table 6: Overall change of LLM performance (F1 score) with different improvement strategies. The performance change is compared with Table 3, where results of $Y^{\prime}$ demonstrate the final counterfactual reasoning outcomes.
Dataset Improve Explicit Variable (§4.2.1) Improve Implicit Variable (§4.2.2) Improve Both
G5 QW3 LM4S GM2.5 G5 QW3 LM4S GM2.5 G5 QW3 LM4S GM2.5
CRASS ++1.8 ++2.3 ++3.5 ++2.0 ++6.2 ++7.4 ++7.9 ++6.8 ++9.0 ++10.1 ++10.5 ++9.3
CLOMO ++2.1 ++3.1 ++4.2 ++2.6 ++5.7 ++6.3 ++6.1 ++6.5 ++9.2 ++10.4 ++11.0 ++9.8
COCO ++1.2 ++1.5 ++1.9 ++1.4 ++4.1 ++5.1 ++5.9 ++5.1 ++5.8 ++6.4 ++6.7 ++6.2
Open-Critic ++1.6 ++2.3 ++2.7 ++2.0 ++3.3 ++8.3 ++11.2 ++4.3 ++5.3 ++11.5 ++14.2 ++5.5
In an end-to-end setting, we incorporate the prior improvements for explicit variable identification (Section 4.2.1), implicit variable reasoning (Section 4.2.2), and their combinatorial strategy to evaluate the overall change in counterfactual reasoning performance. Table 6 presents the results across representative datasets and LLMs.
We have to observations: (i) In general, improving implicit variable reasoning (i.e., for mediators and outcomes) yields more substantial gains in end-to-end performance, as these variables directly influence the final counterfactual predictions (Task IV). In contrast, improvements in explicit variable identification (e.g., for $X,Z,Y$) primarily strengthen the early stages of the pipeline, offering moderate but necessary support. (ii) Notably, the combined strategy achieves the highest overall improvement, although the gains are not strictly additive. This is due to accumulative reasoning errors across the decompositional steps, where inaccuracies in earlier predictions may propagate and compound through the pipeline. These findings highlight both the opportunities and challenges of modular improvements.
@@ Appendix A Complementary Information of Causality and Causal Graph
In evaluation, we rely on a universal prompt template as shown below, wherein only the task-specific contents are replaced for each dataset.
Prompt Template.
**Task Description**
You are asked to perform [Task I / Task II / Task III / Task IV] in decompositional counterfactual reasoning. Follow the definitions of causal variables and causal relations strictly.
**Input Context**
Here is the factual instance from the dataset:
[Insert factual context or multimodal description]
**Intermediate Outputs**
If applicable, use the following ground-truth results from previous tasks:
Exposure (X): [...]
Covariate(s) (Z): [...]
Mediator(s) (M): [...]
Outcome (Y): [...]
**Instruction for the Current Task**
Task I (Variable Identification): [description about domain specific meanings of causal variables X, Z, M, and Y and the identification task]
Task II (Graph Construction): [instruction about constructing the causal graph by listing all directed edges among X, Z, M, Y.
Task III (Intervention Identification): [instruction about identifying which variable is intervened on in the counterfactual query.]
Task IV (Outcome Reasoning): [instruction about inferring] the counterfactual mediator M and outcome Y under the specified intervention.
**Output Format**
Provide the answer using the following structure:
Exposure (X): [...]
Covariate(s) (Z): [...]
Mediator(s) (M): [...]
Outcome (Y): [...]
Causal Edges: [...]
Intervention: [...]
Counterfactual Mediator (M’): [...]
Counterfactual Outcome (Y’): [...]
@@ Appendix B Complementary Experiment
B.1 Setting